OpenStudy (anonymous):

PLEASEEE HELPP ME! f(x) = 9x4 + 5x3 - 4x2 - 7x - 1 what are the exact two inflection points? Kay thanks (:

OpenStudy (amistre64):

derive it twice and solve for 0; that will give you points to test

OpenStudy (amistre64):

also where it is undefined

OpenStudy (anonymous):

36x^3+15x^2-8x-7

OpenStudy (amistre64):

good; thats one derivative :)

OpenStudy (anonymous):

108x^2+30x-8

OpenStudy (amistre64):

nice work; now we either factor it or use the quadratic formula to determine the "zeros"

OpenStudy (anonymous):

This is the part I can never understand at all! The solve for 0 part

OpenStudy (amistre64):

by solving for "0" we are determiningg the places where the slope is a horizontal line

OpenStudy (amistre64):

at those places we tend to have critical values

OpenStudy (amistre64):

http://www.wolframalpha.com/input/?i=108x^2%2B30x-8 to check the roots, wolfram gives us these options

OpenStudy (amistre64):

this version might help in "seeing" what the derivatives look like

OpenStudy (anonymous):

ohh so. 108x^2+30x-8=0 so two values that added together are -8 and multiplied are 30?

OpenStudy (anonymous):

how do I get the answer from that graph :/

OpenStudy (anonymous):

??????

OpenStudy (amistre64):

the zeroes for this might be better found with the quadratic formula such that: -30 +- sqrt(30^2 -4(108)(-8)) x = --------------------------- 2(108) -30 +- sqrt(900 +108(32)) x = --------------------------- 2(108) im just gonna google calculate it .... (-30 + sqrt(900 +108(32)))/(2*108) = 1/6 (-30 - sqrt(900 +108(32)))/(2*108) = -4/9

OpenStudy (amistre64):

so when x=1/6 ; or x=-4/9, we have a critical value to test

OpenStudy (amistre64):

becasue of the nature of the original equation; its a 4th degree, I already know these are the inflection points :)

OpenStudy (anonymous):

Ohhh emmm geee! Thanksssss! You're the best!

OpenStudy (amistre64):

youre welcome